Marla begins walking at 3 mi/h toward the library. Her friend meets her at the half way point and drives her the rest of the way to the library. The distance to the library is 4 miles. How many hours did Marla walk.
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(57344)

Marla begins walking at 3 mi/h toward the library. Her friend meets her at the half way point and drives her the rest of the way to the library. The distance to the library is 4 miles. How many hours did Marla walk.
--------------
Maria Walking DATA:
distance= 2 miles ; rate= 3 mph ; time= d/r = 2/3 hrs
----------
Looks like she walked for 2/3 hr or 40 minutes.
cheers,
Stan H.

Answer by checkley71(8403)

SEEING AS SHE WALKED HALF WAY TO THE LIBRARY SHE WALKED 2 MILES THUS 2=3T OR
T=2/3 OR SHE WALKED 40 MINUTES (2/3 HOUR)